# Perception- and Fidelity-aware Reduced-Reference Super-Resolution Image Quality Assessment

### v1

@@ Abstract
With the advent of image super-resolution (SR) algorithms, how to evaluate the quality of generated SR images has become an urgent task. Although full-reference methods perform well in SR image quality assessment (SR-IQA), their reliance on high-resolution (HR) images limits their practical applicability. Leveraging available reconstruction information as much as possible for SR-IQA, such as low-resolution (LR) images and the scale factors, is a promising way to enhance assessment performance for SR-IQA without HR for reference. In this letter, we attempt to evaluate the perceptual quality and reconstruction fidelity of SR images considering LR images and scale factors. Specifically, we propose a novel dual-branch reduced-reference SR-IQA network, i.e., Perception- and Fidelity-aware SR-IQA (PFIQA). The perception-aware branch evaluates the perceptual quality of SR images by leveraging the merits of global modeling of Vision Transformer (ViT) and local relation of ResNet, and incorporating the scale factor to enable comprehensive visual perception. Meanwhile, the fidelity-aware branch assesses the reconstruction fidelity between LR and SR images through their visual perception. The combination of the two branches substantially aligns with the human visual system, enabling a comprehensive SR image evaluation. Experimental results indicate that our PFIQA outperforms current state-of-the-art models across three widely-used SR-IQA benchmarks. Notably, PFIQA excels in assessing the quality of real-world SR images.
@@ I Introduction
Image super-resolution (SR) technology aims to produce more detailed high-resolution (HR) images from the given low-resolution (LR) images. It has been widely used in various fields, such as security and surveillance[1], medical imaging[2], and remote sensing imaging[3, 4]. Given the remarkable advancements in recent research on SR algorithms, including blind SR[5, 6], lightweight SR models[7, 8], arbitrary scale SR[9], and multimodal SR[10, 11], it has become imperative to evaluate the quality of the generated SR images. This evaluation is crucial for facilitating comparative analysis of reconstruction performance across various SR models and guiding the development of SR algorithms.
@@
Numerous methods have been proposed for image quality assessment (IQA) [12, 13, 14], which can be categorized as subjective or objective. While subjective ones are more reliable, they are impractical due to high costs and external factors. Hence, objective IQA methods that align with subjective evaluations are currently a focus of research. IQT[15], MANIQA[14], AHIQ[16], and MAMIQA[17] are recently proposed IQA methods that achieve excellent subjective consistency. While generic IQA methods have shown satisfactory results, they often neglect the specific characteristics of SR, making them unsuitable to directly apply to SR images. SR algorithms aim to recover detailed information from LR images, so SR-IQA needs to not only focus on the visual quality of the image, but also consider the consistency with the LR image, a factor that generic IQA methods overlook. At the same time, SR images often exhibit mixed degradations, such as blurring, ringing, and aliasing artifacts, which are not effectively addressed by current generic IQA methods.
Figure 1: Overview of proposed PFIQA method. This framework consists of Perception-aware and Fidelity-aware Assessment Branches for SR-IQA.
@@
Recently, there has been an increase in the development of SR-IQA methods. Depending on the availability of lossless reference images, SR-IQA methods can be categorized into full-reference[18], reduced-reference[19, 11], and no-reference methods[20, 21]. For all three SR-IQA paradigms, evaluating the perceptual quality is of paramount importance, as it directly relates to human perceptual judgments and assessment of the given SR images[22, 18]. Moreover, reconstruction fidelity[23, 24] is also highly important since it reflects how faithfully the SR image represents the reference image details and content. Only reference-based SR-IQA methods can effectively consider reconstruction fidelity by comparing the SR image against the reference image. However, in practical scenarios where HR images are unavailable, the application of full-reference SR-IQA methods becomes challenging. Fortunately, the SR task inherently possesses available reference or auxiliary information, such as the paired LR image and the scale factor. These cues have significant reference value for evaluating SR images[18], but are not directly utilized by most SR-IQA methods[21, 20]. Recently, a few reduced-reference SR-IQA methods [19, 18, 25] have adopted CNN-based networks to get feature maps of LR and SR images, and simply calculated the similarity between the feature maps and utilized multi-layer perceptron to regress the quality scores. The entire process can be seen as a pixel-by-pixel comparison between the LR and SR images, which yet neglects the spatial coherence between the LR and SR images, resulting in an insufficient emphasis on reconstruction fidelity.
@@
II Methodology
@@ II-A Feature Extraction
Specifically, a pair of LR and SR images are fed into ViT and ResNet, and then we take out the ViT feature maps of the early stage and ResNet feature maps of different scales. Due to the similarity in the feature extraction process between SR and LR images, we illustrate the process taking the SR image as an example. For ViT, we utilize the outputs from five intermediate blocks $\{0,1,2,3,4\}$ within ViT. The output feature from each block $f_{sr}\in\mathbb{R}^{p\times p\times c}$, where $c=768$, $p=28$, is concatenated into $f_{sr}^{G}\in\mathbb{R}^{p\times p\times 5c}$, and then reshaped into $f_{sr}^{G^{\prime}}\in\mathbb{R}^{5c\times p\times p}$. For ResNet, we extract multi-scale features from four different stages of ResNet, and then interpolate these features to resize them to the same dimensions. After that, we concatenate them along the channel dimension and reshape them to obtain $f_{sr}^{L^{\prime}}\in\mathbb{R}^{3840\times p\times p}$. We apply a convolutional dimensionality reduction to get the final output global features $f_{sr}^{G}\in\mathbb{R}^{256\times p\times p}$ and local features $f_{sr}^{L}\in\mathbb{R}^{256\times p\times p}$. Similarly, the paired LR image undergoes the same process, resulting in the output of global features $f_{lr}^{G}$ and local features $f_{lr}^{L}$.
@@ II-B Adaptive Fusion
II-C Patch-Weighted Quality Regression
@@
TABLE I: Performance comparison on three benchmark datasets. We highlight the best and the second results. Algorithms marked with asterisks (*) have been reproduced, while unmarked ones use results from the original paper.
@@
Generic IQA $\times$ NIQE*[13] 0.4055 0.3916 0.6286 0.5520 0.1452 0.0983 0.3931 0.3473
$\times$ BRISQUE*[12] 0.5568 0.5851 0.5397 0.5085 0.0515 0.0075 0.3827 0.3670
@@
III Experiments
III-A Experimental Settings
@@
Datasets and Evaluation Metrics. The evaluations are implemented on three datasets which are commonly used in the research of SR-IQA, including QADS[30], WIND[31] and RealSRQ[32]. We split the datasets randomly into training and test sets at an 8:2 ratio, repeat the partitioning and evaluation process 5 times for fair comparison, and report the average results as the final performance. We utilize two of the most commonly used metrics, including Pearson Linear Correlation Coefficient (PLCC) and Spearman Rankorder Correlation Coefficient (SRCC).
@@
Implementation Details. We use ViT-B [26] and ResNet50 [27] models pre-trained on ImageNet [36], which are not updated during training. We bilinearly interpolate the LR image to achieve the same resolution as the SR image. For model training, we normalize all input images and randomly crop them into a size of $224\times 224$. Additionally, a random horizontal flip rotation is applied to augment the training data. To optimize the model, we utilize the MSE loss between the predicted score and the ground truth score. During the test phase, we crop the four corners and the center of the original image, and the final score is the average of the scores from these cropped sections. We train PFIQA on a single NVIDIA GeForce RTX3090 GPU. The training process utilizes the AdamW optimizer. The minibatch size is set to 4.
III-B Experimental Results and Analysis
TABLE II: Comparison of leveraging available information of SR algorithms on RealSRQ[32].
@@ III-B Experimental Results and Analysis
[…]
Effects of Reference Information. To further investigate the impact of reference information for SR-IQA, we compare the performance of PFIQA utilizing different reference information or not. In Table II, we can see that: (1) Using only the PA Branch (Table II (a)) can be viewed as a no-reference SR-IQA method, focusing solely on the perceptual quality of the SR image itself, hence leading to sub-optimal performance. (2) Considering the LR image as reference and using only the FA Branch (Table II (b)) can be viewed as a plain reduced-reference SR-IQA method, which accounts for the consistency between the SR and LR images, thereby ensuring reconstruction fidelity to a certain extent. Compared to Table II (a), this brings a significant performance improvement. (3) Combining the PA Branch and FA Branch (Table II (c)), integrates both perceptual quality and reconstruction fidelity, thus obtaining better performance compared to Table II (a,b). (4) Building upon considering the LR image as the reference, and additionally taking into account the scale factor (Table II (d)), PFIQA maximally leverages the available information from SR algorithms, thereby achieving the best SR-IQA performance.
TABLE III: Comparison of different visual feature extraction and fusion strategies on RealSRQ[32].
@@
Effects of Different Visual Feature Extraction Methods and Fusion Strategies. Since we use ViT and ResNet to extract global and local visual features from images, it is necessary to investigate how they impact the performance of SR-IQA. In Table III, we can find that: (1) Using only each of the visual features extracted by ViT and ResNet (Table III (a,b)) can achieve satisfactory SR-IQA results compared to other SOTA methods in Table I. (2) Utilizing both global and local visual features (Table III (c)), leads to better performance compared to Table III (a,b). This demonstrates the complementary nature of global and local features for comprehensive SR-IQA. (3) Further analysis of the adaptive fusion of global and local features, as indicated in Table III (d), reveals that PFIQA attains superior performance. This demonstrates the efficacy of the AFMs in learning the importance of features from global and local features.
IV Conclusion
In this letter, we propose a novel reduced-reference Perception- and Fidelity-aware SR-IQA (PFIQA) network, which integrates LR images and the scale factors as prior knowledge to comprehensively assess the perceptual quality and reconstruction fidelity of SR images. We leverage the merits of global modeling of ViT and local relation of CNN to enable comprehensive visual feature extraction, and incorporate the scale factors to obtain features that are more relevant to SR-IQA. The extensive results from three benchmark datasets showcase the efficacy of PFIQA.

### v2

@@ Abstract
With the advent of image super-resolution (SR) algorithms, how to evaluate the quality of generated SR images has become an urgent task. Although full-reference methods perform well in SR image quality assessment (SR-IQA), their reliance on high-resolution (HR) images limits their practical applicability. Leveraging available reconstruction information as much as possible for SR-IQA, such as low-resolution (LR) images and the scale factors, is a promising way to enhance assessment performance for SR-IQA without HR for reference. In this paper, we attempt to evaluate the perceptual quality and reconstruction fidelity of SR images considering LR images and scale factors. Specifically, we propose a novel dual-branch reduced-reference SR-IQA network, i.e., Perception- and Fidelity-aware SR-IQA (PFIQA). The perception-aware branch evaluates the perceptual quality of SR images by leveraging the merits of global modeling of Vision Transformer (ViT) and local relation of ResNet, and incorporating the scale factor to enable comprehensive visual perception. Meanwhile, the fidelity-aware branch assesses the reconstruction fidelity between LR and SR images through their visual perception. The combination of the two branches substantially aligns with the human visual system, enabling a comprehensive SR image evaluation. Experimental results indicate that our PFIQA outperforms current state-of-the-art models across three widely-used SR-IQA benchmarks. Notably, PFIQA excels in assessing the quality of real-world SR images.
@@ I Introduction
Image super-resolution (SR) technology aims to produce more detailed high-resolution (HR) images from the given low-resolution (LR) images. It has been widely used in various fields, such as security and surveillance[1], medical imaging[2], and remote sensing imaging[3, 4]. Given the remarkable advancements in recent research on SR algorithms, including blind SR[5, 6], lightweight SR models[7, 8, 9], arbitrary scale SR[10], and multimodal SR[11, 12], it has become imperative to evaluate the quality of the generated SR images. This evaluation is crucial for facilitating comparative analysios of reconstruction performance across various SR models and guiding the development of SR algorithms.
@@
Numerous methods have been proposed for image quality assessment (IQA) [13, 14, 15, 16, 17], which can be categorized as subjective or objective. While subjective ones are more reliable, they are impractical due to high costs and external factors. Hence, objective IQA methods that align with subjective evaluations are currently a focus of research. IQT[18], MANIQA[17] and AHIQ[19] are recently proposed IQA methods that achieve excellent subjective consistency. While generic IQA methods have shown satisfactory results, they often neglect the specific characteristics of SR, making them unsuitable to directly apply to SR images. SR algorithms aim to recover detailed information from LR images, so SR-IQA needs to not only focus on the visual quality of the image, but also consider the consistency with the LR image, a factor that generic IQA methods overlook. At the same time, SR images often exhibit mixed degradations, such as blurring, ringing, and aliasing artifacts, which are not effectively addressed by current generic IQA methods.
@@
Recently, there has been an increase in the development of SR-IQA methods[20]. Depending on the availability of lossless reference images, SR-IQA methods can be categorized into full-reference[21], reduced-reference[22, 12], and no-reference methods[23, 24]. For all three SR-IQA paradigms, evaluating the perceptual quality is of paramount importance, as it directly relates to human perceptual judgments and assessment of the given SR images[25, 21]. Moreover, reconstruction fidelity[26, 27] is also highly important since it reflects how faithfully the SR image represents the reference image details and content. Only reference-based SR-IQA methods can effectively consider reconstruction fidelity by comparing the SR image against the reference image. However, in practical scenarios where HR images are unavailable, the application of full-reference SR-IQA methods becomes challenging. Fortunately, the SR task inherently possesses available reference or auxiliary information, such as the paired LR image and the scale factor. These cues have significant reference value for evaluating SR images[21], but are not directly utilized by most SR-IQA methods[24, 23]. Recently, a few reduced-reference SR-IQA methods [22, 21, 28] have adopted CNN-based networks to get feature maps of LR and SR images, and simply calculated the similarity between the feature maps and utilized multi-layer perceptron to regress the quality scores. The entire process can be seen as a pixel-by-pixel comparison between the LR and SR images, which yet neglects the spatial coherence between the LR and SR images, resulting in an insufficient emphasis on reconstruction fidelity.
@@
II Related work
II-A Image Quality Assessment
We briefly review generic image quality assessment (IQA) methods, which can be broadly categorized into full-reference IQA (FR-IQA), no-reference IQA (NR-IQA) and reduced-reference IQA (RR-IQA).
For FR-IQA, the most widely used metrics are PSNR and SSIM[29]. However, despite their simplicity and ease of optimization, these metrics correlate poorly with HVS[19]. To address the limitations of conventional IQA methods, various learning-based FR-IQA approaches have been recently proposed[18, 19]. IQT[18] utilizes an encoder-decoder transformer with trainable quality embeddings for superior performance. AHIQ[19] predicts image quality at the patch level, excelling in GAN-generated distortions. For NR-IQA, current methods emphasize generality and can be further categorized into two types: natural scene statistics (NSS)-based metrics[16, 15] and learning-based metrics[30, 17]. NSS-based NR-IQA methods require manual extraction of image features, and these hand-crafted feature representations are often not effective. Subsequently, learning-based methods have demonstrated superior performance. CNNIQA[30] introduces a pioneering CNN for IQA, directly learning from image pixels without the need for hand-crafted features. MANIQA[17] proposes a multi-dimensional attention mechanism for interactions in both channel and spatial domains, employing ViT and novel modules to enhance global and local interactions. For RR-IQA, CKDN[31] learns a reference space from degraded images to extract knowledge from pristine-quality images.
Although current IQA methods demonstrate satisfactory results, they often fail to account for the specific characteristics of SR images, making them inadequate for direct application in SR-IQA. SR images typically exhibit various mixed degradations, including blurring, ringing, and aliasing artifacts, which current generic IQA methods are not effective in addressing these issues. Therefore, there is a pressing need for specialized SR-IQA methods that take into consideration the unique degradation patterns and characteristics of SR images.
II-B Super-Resolution Image Quality Assessment
SR-IQA approaches can be divided into two main categories: hand-crafted feature-based methods[32, 33, 34, 35] and learning-based methods[22, 24, 23, 21].
The hand-crafted feature-based methods perform regression to predict the quality scores of SR images. Yeganeh et al.[32] employ three statistical features for reduced-reference SR-IQA, including sub-image frequency energy falloff statistics, sub-image local dominant orientation statistics, and spatial continuity statistics. NRQM[33]adopts three types of low-level statistical features in both spatial and frequency domains to quantify SR artifacts for no-reference SR-IQA. SIS[34] leverages structure-texture decomposition to evaluate the visual quality of SR images. However, these types of methods have shown limited effectiveness in practice.
For learning-based methods, the goal is to automatically map images to their corresponding quality scores. Li et al.[24] introduce C2MT, a class-aware multi-task transformer that uses supervised contrastive learning and active learning. PCST[21] addresses the challenge of balancing referenced and no-reference scores by integrating multi-scale and saliency information into deep learning modules. However, in real-world scenarios where HR images are unavailable, full-reference SR-IQA methods become impractical. DISQ[22], inspired by the VGG network, employs a dual-stream deep neural network architecture by utilizing LR images, which is considered a reduced-reference SR-IQA method. However, in addition to LR images, SR methods can also leverage certain priors, such as the scale factor. Fu et al.[23] present a scale-guided hypernetwork framework (SGH) for no-reference SR-IQA, demonstrating improved performance over existing IQA metrics through scale-adaptive evaluation. While these cues are valuable for SR-IQA, most SR-IQA methods do not directly utilize them[24, 23]. Concurrently, some recent reduced-reference approaches[22, 21, 28] use CNNs to extract and compare LR and SR image features, then employ MLPs for quality scoring. However, this pixel-level comparison fails to account for spatial coherence, undermining the assessment of reconstruction fidelity. In this work, our proposed PFIQA utilizes LR images and scale factors as priors, combines ViT and CNN for comprehensive feature extraction, and adopts fine-grained patch-wise prediction for SR-IQA to assess both the perceptual quality and reconstruction fidelity of SR images.
Figure 1: Overview of proposed PFIQA method. This framework consists of Perception-aware Assessment Branch (PA Branch) and Fidelity-aware Assessment Branch (FA Branch) for SR-IQA.
III Methodology
@@ III-A Feature Extraction
Specifically, a pair of LR and SR images are fed into ViT and ResNet, and then we take out the ViT feature maps of the selected stages and ResNet feature maps of different scales. Due to the similarity in the feature extraction process between SR and LR images, we illustrate the process taking the SR image as an example. For ViT, we utilize the outputs from five stages within the ViT backbone. The output feature from each block $f_{sr}\in\mathbb{R}^{p\times p\times c}$, where $c=768$, $p=28$, is concatenated into $f_{sr}^{G}\in\mathbb{R}^{p\times p\times 5c}$, and then reshaped into $f_{sr}^{G^{\prime}}\in\mathbb{R}^{5c\times p\times p}$. For ResNet, we extract multi-scale features from four different stages of ResNet, and then interpolate these features to resize them to the same dimensions. After that, we concatenate them along the channel dimension and reshape them to obtain $f_{sr}^{L^{\prime}}\in\mathbb{R}^{3840\times p\times p}$. We apply a convolutional dimensionality reduction to get the final output global features $f_{sr}^{G}\in\mathbb{R}^{256\times p\times p}$ and local features $f_{sr}^{L}\in\mathbb{R}^{256\times p\times p}$. Similarly, the paired LR image undergoes the same process, resulting in the output of global features $f_{lr}^{G}$ and local features $f_{lr}^{L}$.
@@ III-B Adaptive Fusion
TABLE I: An overview of widely-used SR-IQA benchmarks.
Datasets LR Images SR Images Number of SR methods Scaling factors Labels Synthetic / Realistic
[…]
@@
III-C Patch-Weighted Quality Regression
@@
IV Experiments
TABLE II: Performance comparison on three benchmark datasets. We highlight the best and the second results. Algorithms marked with asterisks (*) have been reproduced, while unmarked ones use results from the original paper.
@@ IV Experiments
Generic IQA HR PSNR* 0.3024 0.2949 0.7455 0.7431 0.0812 0.0951 0.3764 0.3777
HR SSIM*[29] 0.4862 0.4833 0.2899 0.4216 0.1062 0.1319 0.2941 0.3456
$\times$ NIQE*[16] 0.3043 0.3478 0.2393 0.3195 0.0268 0.0981 0.1901 0.2551
$\times$ BRISQUE*[15] 0.5282 0.5449 0.7475 0.7552 0.0515 0.0076 0.4424 0.4359
@@
Figure 2: Examples of predicted scores on RealSRQ[40] include DISQ[22] (Previous SR-IQA SOTA) / PFIQA (Ours) / MOS. We consistently crop and enlarge each image for better visualization. The images are arranged in order of increasing image quality, with corresponding MOS values progressively increasing.
IV-A Experimental Settings
@@
Datasets and Evaluation Metrics. The evaluations are implemented on three datasets which are commonly used in the research of SR-IQA, including QADS[34], WIND[32] and RealSRQ[40]. Table I describes the details of the above datasets. We split the datasets randomly into training and test sets at an 8:2 ratio, repeat the partitioning and evaluation process 5 times for fair comparison, and report the average results as the final performance. We utilize two of the most commonly used metrics, including Pearson Linear Correlation Coefficient (PLCC) and Spearman Rank-order Correlation Coefficient (SRCC). PLCC assesses the linear correlation between mean opinion scores (MOSs) and predicted quality scores. It is used to measure the accuracy of IQA algorithm predictions, while SRCC measures the monotonicity of IQA algorithm predictions. Higher values of these two metrics indicate better performance of the IQA method.
@@
Implementation Details. We use ViT-B/8 [36] and ResNet-50 [37] models pre-trained on ImageNet [41], which are not updated during training. We bilinearly interpolate the LR image to achieve the same resolution as the SR image. For model training, we normalize all input images and randomly crop them into a size of $224\times 224$. Additionally, a random horizontal flip rotation is applied to augment the training data. To optimize the model, we utilize L2 Loss between the predicted score and the MOS, and the batch size is 4. During the test phase, we crop the four corners and the center of the original image, and the final score is the average of the scores from these cropped sections. We train PFIQA on a single NVIDIA GeForce RTX3090 GPU. The training process utilizes the AdamW optimizer, setting the maximum number of training epochs to 200.
IV-B Main Results
Figure 3: Scatter plots of ground-truth mean opinion scores (MOSs) against predicted scores of six generic IQA methods, two SR-IQA methods and proposed PFIQA on RealSRQ[40] dataset. The MOS values have been scaled to range between 0 and 1. Blue points represent the results of the corresponding methods, and the linear fitting of all points is marked by a red straight line.
[…]
Qualitative Analysis. Fig. 2 presents examples of MOS values and predicted scores from DISQ and our PFIQA methods. We can observe that the scores predicted by PFIQA demonstrate better consistency with MOS values compared to DISQ. Furthermore, predicted scores from PFIQA show a strong alignment with the HVS.
To further validate the effectiveness of the proposed PFIQA, Fig. 3 presents scatter plots comparing nine metrics with predicted scores through linear fitting. The vertical axis represents subjective MOS values, while the horizontal axis corresponds to predicted scores. Examining the resulting scatter plots, we observe that the points plotted by PFIQA are more compactly distributed and closer to the diagonal line compared to other methods. These impressive results indicate that the quality scores predicted by PFIQA demonstrate a more consistent correlation with subjective assessment of SR images than other comparative methods.
IV-C Ablation Study
In this section, we conduct extensive ablation experiments to investigate the effects of various factors on PFIQA for SR-IQA. These experiments include utilizing different reference information, feature extraction and fusion strategies, and fine-tuning techniques. All experiments in this section are performed on the RealSRQ [40] dataset.
Effects of Reference Information. To further investigate the impact of reference information for SR-IQA, we compare the performance of PFIQA utilizing different reference information or not. In Table III, we can see that: (1) Using only the PA Branch (Table III (a)) can be viewed as a no-reference SR-IQA method, focusing solely on the perceptual quality of the SR image itself, hence leading to sub-optimal performance. (2) Considering the LR image as reference and using only the FA Branch (Table III (b)) can be viewed as a plain reduced-reference SR-IQA method, which accounts for the consistency between the SR and LR images, thereby ensuring reconstruction fidelity to a certain extent. Compared to Table III (a), this brings a significant performance improvement. (3) Combining the PA Branch and FA Branch (Table III (c)), integrates both perceptual quality and reconstruction fidelity, thus obtaining better performance compared to Table III (a,b). (4) Building upon considering the LR image as the reference, and additionally taking into account the scale factor (Table III (d)), PFIQA maximally leverages the available information from SR algorithms, thereby achieving the best SR-IQA performance.
TABLE III: Comparison of leveraging available information of SR algorithms on RealSRQ[40].
@@ IV-C Ablation Study
TABLE IV: Comparison of different visual feature extraction and fusion strategies on RealSRQ[40].
@@
TABLE V: Comparison of using different different fine-tuning strategies on RealSRQ[40].
# ViT ResNet PLCC SRCC
Updated Updated
(a) 0.9025 0.8277
(b) 0.9195 0.8539
(c) 0.9065 0.8305
(d) 0.9269 0.8597
Effects of Different Feature Extraction and Fusion Strategies. Since we use ViT and ResNet to extract global and local visual features from images, it is necessary to investigate how they impact the performance of SR-IQA. In Table IV, we can find that: (1) Using only each of the visual features extracted by ViT and ResNet (Table IV (a,b)) can achieve satisfactory SR-IQA results compared to other methods in Table II. (2) Utilizing both global and local visual features (Table IV (c)), leads to better performance compared to Table IV (a,b). This demonstrates the complementary nature of global and local features for comprehensive SR-IQA. (3) Further analysis of the adaptive fusion of global and local features, as indicated in Table IV (d), reveals that PFIQA attains superior performance. This demonstrates the efficacy of the AFMs in learning the importance of features from global and local features.
Effects of Different Fine-tuning Strategies. To investigate the impact of fine-tuning ViT and ResNet on the performance of SR-IQA, we conducted experiments with different parameter updating strategies. In Table V, we observe that: (1) Keeping both pre-trained ViT and ResNet parameters fixed (Table V (d)) yields the best performance, with PLCC of 0.9269 and SRCC of 0.8597. This suggests that the pre-trained knowledge from these models is highly valuable for SR-IQA task. This approach helps maintain the model’s generalization capability and mitigates the risk of overfitting on small, specialized datasets like those used in SR-IQA tasks. (2) Updating either ViT or ResNet individually (Table V (a,b)) leads to a decrease in performance compared to the non-updated version (Table V (d)). This indicates that fine-tuning on the small SR-IQA dataset may lead to overfitting and loss of generalized features learned during pre-training. (3) Updating both ViT and ResNet (Table V (c)) results in further performance degradation. This reinforces the observation that preserving the pre-trained knowledge is crucial for the tsak of SR-IQA task, especially given the limited scale of the SR-IQA datasets.
V Conclusion
In this paper, we propose a novel reduced-reference Perception- and Fidelity-aware SR-IQA (PFIQA) network, which integrates LR images and the scale factors as prior knowledge to comprehensively assess the perceptual quality and reconstruction fidelity of SR images. We leverage the merits of global modeling of ViT and local relation of CNN to enable comprehensive visual feature extraction, and incorporate the scale factors to obtain features that are more relevant to SR-IQA. The extensive results from three benchmark datasets showcase the efficacy of PFIQA, and the evaluation results are highly consistent with HVS.